OpenStudy (anonymous):

write the standard equation of a circle with center (-4 0) and radius 3

OpenStudy (solomonzelman):

A standard equation of a circle with center \(\large\color{black}{ (h,k) }\) and radius \(\large\color{black}{ r }\) is: \(\large\color{black}{ (x-h)^2+(y-k)^2=r^2 }\)

OpenStudy (solomonzelman):

A couple of examples: ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ example 1: A standard equation of a circle with center \(\large\color{black}{ (1,2) }\) and radius \(\large\color{black}{3 }\) is: \(\large\color{black}{ (x-1)^2+(y-2)^2=9 }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ example 2: A standard equation of a circle with center \(\large\color{black}{ (3,0) }\) and radius \(\large\color{black}{1 }\) is: \(\large\color{black}{ (x-3)^2+y^2=1 }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ example 3: A standard equation of a circle with center \(\large\color{black}{ (0,-4) }\) and radius \(\large\color{black}{5 }\) is: \(\large\color{black}{ x^2+(y+4)^2=25 }\)

OpenStudy (solomonzelman):

if the examples don't help let me know.... basically, if you have to say anything, you are more than welcome to do so....

OpenStudy (anonymous):

A: (x-4)^2 +y^2=3 B:(x+4)^2 +y^2=3 C:(x-4)^2 +y^2=9 D:(x+4)^2 +y^2=9

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I see the options, so can you go ahead and complete the equation? ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ A standard equation of a circle with center \(\large\color{black}{ (-4,0) }\) and radius \(\large\color{black}{3 }\) is: \(\large\color{black}{ (x~~~~~~~)^2+(y~~~~~~~)^2=~~~~.}\)

OpenStudy (solomonzelman):

can you feel in gaps, based on the examples that I went through

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (alexandervonhumboldt2):

D

OpenStudy (anonymous):

can you help with a few more @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes post in separate post and tag me